40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 16:46:31 2019, gautam, Update, IOO, IMC not locking PCdrive_RMS.png
    Reply  Tue Jun 4 10:56:31 2019, gautam, Update, IOO, IMC diagnostics IMCdiag.pdf
       Reply  Thu Jun 6 22:11:53 2019, Koji, Update, IOO, IMC diagnostics 
          Reply  Thu Jun 13 18:01:18 2019, aaron, Update, IOO, IMC diagnostics 
             Reply  Fri Jun 14 13:10:00 2019, aaron, Update, IOO, IMC diagnostics 
                Reply  Mon Jun 17 12:37:16 2019, aaron, Update, IOO, IMC diagnostics 
                   Reply  Tue Jun 18 20:35:07 2019, aaron, Update, IOO, IMC diagnostics 
                      Reply  Wed Jun 19 19:12:51 2019, aaron, Update, IOO, IMC diagnostics PMC_Error_Spectrum.pdfPDH_signal.pdfPDH_signal_full.pdf
                         Reply  Mon Jun 24 15:49:05 2019, aaron, Update, IOO, IMC diagnostics sweep_config_updated.png
                            Reply  Tue Jul 2 19:36:18 2019, aaron, Update, IOO, IMC diagnostics PDH_error.pdfPMC_Error_Spectrum.pdf
          Reply  Fri Jun 21 19:36:26 2019, Koji, Update, IOO, IMC diagnostics IMC_error.pdf
             Reply  Sun Jun 23 08:09:53 2019, gautam, Update, IOO, NPRO diagnostics consolidatedOscopeScreenCaps.pdfRINcomp.pdfPLL_OLTF.pdfPLLnoise.pdf
                Reply  Sun Jun 23 09:36:32 2019, gautam, Update, IOO, IMC is locking normally again PMCdiag.pdf
                   Reply  Sun Jun 23 14:43:14 2019, Koji, Update, IOO, IMC is locking normally again 
                   Reply  Mon Jun 24 08:12:10 2019, gautam, Update, IOO, IMC is locking normally again IMCdutycycle.png
                      Reply  Mon Jun 24 11:48:35 2019, gautam, Update, IOO, IMC in-loop error spectra and OLTF IMCfreqNoise.pdfIMC_OLTF.pdfIMC_CMboard.jpg
Message ID: 14721     Entry time: Tue Jul 2 19:36:18 2019     In reply to: 14693
Author: aaron 
Type: Update 
Category: IOO 
Subject: IMC diagnostics 

The latest in my fling with the PMC. Though PMC trans is back to nominal levels (~0.713 V), we'd still like to understand the PMC noise.

Last time, I took some spectra with the RF probe (Agilent 41800A). I had already measured the PDH error signal by sweeping the PZT at ~1 Hz. The notebook I used for analysis has been updated in /users/aaron/analysis/PDH_calibrate.ipynb. The analysis was the following:

  • fit the PDH error signal, assuming a 35.5MHz modulation frequency. Here are the (approximate) fit parameters:
    • Mapping of PZT mon voltage to Hz: 5.92 Hz/V_{PZT_mon}
    • P_carrier*P_signal: 0.193 W^2
    • HV mon voltage on resonance: 0.910 V
    • Error signal far off resonance: 0.249 V
    • Transmission: 0.00238
      • ​yikes. The nominal transmission is T=0.003. I let this parameter be free as a check, and to avoid overconstraining the data; is this consistent with measurements of the PMC optics' transmission?
    • Length: 0.0210 m
      • This is consistent with the nominal PMC length
  • Using the fit of the full PDH error signal, I am able to plot error signal vs frequency, and fit the linear portion of the carrier PDH signal. The results of this fit are:
    • -9.75e-7 V_PDH per Hz
    • 0.105 V error signal at DC
  • I then divide the power spectra by the squared slope of the linear fit above (V_PDH^2/Hz^2) to get the spectra in Hz^2/rtHz
    • I've plotted both the spectrum I took directly at the mixer I using the agilent probe, as well as the spectrum taken by sending the PMC servo card's mixer mon to an SR560 (G=2) then to the spectrum analyzer

There are a few problems remaining:

  • There should be a gain of 100 between the mixer I and the servo board's mixer out. It's not clear to me that this is reflected in the spectra. Moreover, the header files on the spectra I grabbed from the Agilent say that the R (mixer I) channel has 20dB of input attenuation, which is also not reflected. If I have swapped the two spectra and not accounted for either the gain of the servo card or the attenuation of the spectrum analyzer, these two gains would cancel, but I'm not confident that's what's going on.
Attachment 1: PDH_error.pdf  42 kB  Uploaded Tue Jul 2 21:16:37 2019  | Hide | Hide all
PDH_error.pdf
Attachment 2: PMC_Error_Spectrum.pdf  94 kB  Uploaded Tue Jul 2 21:16:49 2019  | Hide | Hide all
PMC_Error_Spectrum.pdf
ELOG V3.1.3-